OpenStudy (anonymous):

Write the complex number in polar form with argument θ between 0 and 2π. −2 + 2i

OpenStudy (zzr0ck3r):

for \(c=a+bi \ \in \mathbb{C} \) radius is given by \(r = \sqrt{a^2+b^2}\) we know that your point lies in the 2nd quadrant in the complex plane so \(\theta=\arctan(\frac{2}{-2})+\pi=\arctan(-1)+\pi\) in general if the point is in the 1st or 3rd quadrant we have \(a+bi\iff(r,\theta)=(\sqrt{a^2+b^2},\arctan(\frac{b}{a}))\) else \[a+bi\iff(r,\theta)=(\sqrt{a^2+b^2},\arctan(\frac{b}{a})+\pi))\]

OpenStudy (zzr0ck3r):

that should say if the point is in the 1st or \(\color{red}{4th}\) quadrant

OpenStudy (anonymous):

thank you

OpenStudy (zzr0ck3r):

no problem